OpenStudy (anonymous):

Simplify the expression where possible. (6 ^3) ^-3

OpenStudy (anonymous):

Remember: \[a^{-b} = \frac{1}{a^b}\]

OpenStudy (anonymous):

So \[(6^3)^{-3} = \frac{1}{(6^3)^3}\]